Player lI C D Player A 6,3 2,6 В 4,3 8,1 Suppose that in the game above Player 2 plays strategy C with probability q:03. The value for ET71(A, q) is: a. 8.3 b. 3.2 C.3.9 d. 4.6
Answer #1

In the question, it is given that player 2 will choose strategy C with probability 0.3

So the probability of strategy D chosen by player 2 will be =(1 - 0.3) = 0.7

In the question , we are asked to find out the expected profit of player 1 if he chooses strategy A.

Player 1 will have a profit of 6 when he chooses strategy A and player 2 chooses strategy C. And we know that probability of player 2 choosing strategy C is 0.3.

Player 1 will have a profit of 2 when he chooses strategy A and player 2 chooses strategy D. And we know that probability of player 2 choosing strategy D is 0.7.

Expected profit = Profit * probability of having profit

So, expected profit = 6*0.3 + 2*0.7 = 1.8 + 1.4 = 3.2

So option B is the correct answer.